OpenStudy (anonymous):

You flipped two coins. you got two heads 28 times, heads and tails 22 times, tails and heads 34 times, and two tails 16 times. how does the experimental probability of getting two heads compare to the theoretical probability of getting two heads? A) Experimental is 3% greater than theoretical. B) Theoretical is 3% greater than experimental. C) Experimental is equal to theoretical. B) Experimental is 7% greater than theoretical.

OpenStudy (campbell_st):

well there were 100 tosses experimental probability for P(H, H) = 28/100 theoretical probability = 1/4 so convert both fractions to percentages and compare them to your answer choices

OpenStudy (anonymous):

thx! that would have taken me forever to do!